The train, with a mass of 360 t, after leaving the station gains a speed of 36km / h. Find the work done by the locomotive during the first minute. Force It is immutable


1
Expert's answer
2021-03-31T09:31:54-0400

Let's assume that the locomotive starts from rest. Then, we can find the distance that the locomotive travels during 1 minute from the kinematic equation:


"d=\dfrac{1}{2}(v_0+v)t,""d=\dfrac{1}{2}\cdot(0+10\ \dfrac{m}{s})\cdot60\ s=300\ m."

We can find the acceleration of the locomotive as follows:


"a=\dfrac{v-v_0}{t}=\dfrac{10\ \dfrac{m}{s}}{60\ s}=0.17\ \dfrac{m}{s^2}."

Then, we can find the force that acts on the locomotive:


"F=ma=360\cdot10^3\ kg\cdot0.17\ \dfrac{m}{s^2}=61200\ N."

Finally, we can find the work done by the locomotive during the first minute:


"W=Fd=61200\ N\cdot300\ m=1.84\cdot10^7\ J."